OpenStudy (anonymous):

Hi, i ask this the other day and if someone can solve it with answer and explains, will fan and medal: When a beam makes an angle of 40 degrees with the ground, the top of the beam is 8 feet above the ground. There a telephone wires near by and the worker worried that the beam may hit the wires. When the beam makes a angle of 60 degrees with the ground, the wires are 2 feet above the beam. Will the beam hit the ground if the crew continues to raise it?

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

I almost have the answers, i think?! call L is the hypotenuse of the figure: Sin 40 degree=8ft/L L=8ft/(Sin^ 40 )≈ 12.45 Call h is the lenght from thr beam to the ground at angle 60 degre: Sin 60=h/12.45 h=12.45(Sin 60) ≈ 10,78ft

OpenStudy (danica518):